You are the IT Project Manager of a company and there are two projects being considered for selection. Project-I: Developing a website is being considered against Project-II: Developing a Library software. The initial project costs and other details are given below:

 

Project-I: , , CF: 5,000, 10,000, 10,000, 3,000, 2,000, Salvage Value=1,000,

Project-II: , , CF: 20,000, 10,000, 5,000, 3,000, 2,000, Salvage Value=2,000,

Using the NPV, which project should be selected and why?


1
Expert's answer
2021-11-16T18:36:41-0500

Net Present Value, NPV= "\displaystyle\sum_{t=0}^n \frac{R_t}{(1+i)^t}"

t = Time period

Rt = Cash Flow in the time period

i = Interest rate or Discount rate

We need to subtract the initial Capital invested to calculate the true NPV

As given in the problem,

Case - I

Initial Investment = 20000

t = 5

CFt or Rt = 5000, 10000, 10000, 3000, 2000

i = 10%

NPV = CF1/(1+.1)^5 + CF1/(1+.1)^5 + CF1/(1+.1)^5 + CF1/(1+.1)^5 + CF1/(1+.1)^5 - Initial Investment

NPV = 5000/(1.1^1) + 10000/(1.1^2) + 10000/(1.1^3) + 3000/(1.1^4) + 2000/(1.1^5) - 20000

= 3613.94

Case - II

Initial Investment = 30000

t = 5

CFt or Rt = 20000, 10000, 5000, 3000, 2000

i = 10%

NPV = CF1/(1+.1)^5 + CF1/(1+.1)^5 + CF1/(1+.1)^5 + CF1/(1+.1)^5 + CF1/(1+.1)^5 - Initial Investment

NPV = 20000/(1.1^1) + 10000/(1.1^2) + 5000/(1.1^3) + 3000/(1.1^4) + 2000/(1.1^5) - 30000

= 3493.73

As we can see that the NPV for first project is better than second project. Hence, pursuing first project is better. This is considering no Salvage value in the project. Normally, Salvage value is not considered to be part of Cash Flow and hence not used in calculation of Net Present Value.